# Hybrid Higher-Order Topological Skin Modes in the Two-Dimensional Su–Schrieffer–Heeger Model with Nonreciprocal Hoppings

### v1

@@ Abstract
The coexistence of edge states and skin effects provides the topologically protected localized states at the corners of two-dimensional systems. In this paper, we realize such corner states in the two-dimensional Su–Schrieffer–Heeger model with the nonreciprocal hoppings. For the characterization of the real line gap topology, we introduce the $\mathbb{Z}_{4}$ Berry phase protected by generalized four-fold rotational symmetry. From the physical picture of the adiabatic connection, we find that the value of the $\mathbb{Z}_{4}$ Berry predicts the position of edge states. Additionally, by using the winding number, we characterize the point gap topology of the edge spectra. From the results of these characterizations by the first-order topological invariants, we find that the pair of values of the $\mathbb{Z}_{4}$ Berry phase and the winging number yields the position of the topologically protected localized states.
@@
Introduction: Topological materials which are characterized by the topology in the bulk have been actively pursued [1, 2, 3, 4, 5, 6, 7, 8]. Additionally, the description of these materials are enriched by symmetry [9, 10, 11, 12, 13, 14, 15]. For instance, the presence of time-reversal symmetry provides the $\mathbb{Z}_{2}$ topological properties [9]. Moreover, it turns out that crystalline symmetry plays an key role for the determination of the band topology [16, 17, 18, 19, 20, 21, 22, 23, 25, 24, 26]. For example, the coexistence of the time-reversal symmetry and inversion symmetry enables us to introduce the form of $\mathbb{Z}_{2}$ indicator determined from the parity at the high-symmetry points in the two-dimensional Brillouin zone [17].
@@
In the past decades, the notion of the topological protection has been extended to a new area: the non-Hermitian systems [27, 28, 29, 30, 31, 32]. Since the eigenvalues may become complex, many exotic topological phenomena have been provided [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], such as the skin effects [47, 48, 49, 50, 51, 52, 53, 54, 55, 56] and exceptional points (rings) [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. Particularly, a distinct difference from the case in Hermitian systems is the emergence of two types of the gap, i.e. a line gap and a point gap [44]. While the line gap topology can be understood by mapping onto the Hermitian cases, the nontrivial point gap topology provides novel phenomena without analogies in the Hermitian case, such as the anomalous sensitivity of the energy spectrum to the presence or absence of boundaries, i.e., the skin effect. In addition, as is the case for the Hermitian systems, these unique topological phenomena are enriched by symmetry [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84]. These phenomena have been reported in a wide range of systems such as open quantum systems [58, 85, 86], electrical circuits [87, 88, 89, 74, 90, 91], phononic systems [92, 93, 94], and photonic crystals [95, 96, 97, 98], and so on.
@@
Along with the above developments, a novel class of topological insulators has been extensively investigated, which is called higher-order topological insulators. In $d$-dimensional systems, while the conventional topological insulators host $(d-1)$-dimensional boundary states, higher-order topological insulators host $(d-2)$- or fewer dimensional boundary states [99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 113, 112, 114, 115, 116]. For the characterization of the topological phase and the prediction of the emergence of these unique boundary states, many efforts have been devoted. Remarkably, the presence of crystalline symmetry plays an important role for the description of the higher-order topology [99, 100, 101, 102, 103, 107, 108, 109, 110, 111, 113, 112, 114, 115, 116].
@@
Recently, the notion of the higher-order topology has also been extended to non-Hermitian systems [117, 118, 119, 120]. In particular, the coexistence of edge states and the skin effect in two-dimensional systems provides the topologically protected corner states, which are called the hybrid higher-order topological skin modes [121, 122, 123, 124, 125, 126, 127]. These novel states in non-Hermitian systems represent a kind of the higher-order skin effect [128, 129, 130]. Furthermore, these unique localized states are characterized by the first- and second-order topological invariants such as the Chern number [121, 124], the Wess-Sumino term [128], and the symmetry indicator [129, 77].
@@
In this paper, we demonstrate the realization of the hybrid higher-order topological skin modes in the two-dimensional Su-Schrieffer-Heeger (SSH) model with the nonreciprocal hoppings. Specifically, we characterize the first-order topology by using $\mathbb{Z}_{4}$ Berry phase protected by generalized four-fold rotational symmetry and the winding number on the cylinder geometry. From the nontrivial value of the $\mathbb{Z}_{4}$ Berry phase, we show that there exists the bulk-edge correspondence between this topological invariant and edge states. Additionally, by using the winding number, we realize the skin effect arising from the energy spectra of edge states. Combining the $\mathbb{Z}_{4}$ Berry phase and the nontrivial point gap topology of these edge spectra, we demonstrate that the pair of these topological invariants indicates the position of the localized corner in two dimension.
@@
Fig. 1: (Color online) (a) Schematic figure of the two-dimensional SSH model with the nonreciprocal hoppings. The unit cell is enclosed by the black dashed box. The cross denotes the center of the generalized four-fold rotation. (b), (c), and (d) show the energy spectrum for $(\delta t,\delta\gamma)=(-0.7,0.1)$, $(0,0.1)$ and $(0.7,0.1)$. Red lines denote the real line gaps where we focus. (e) Sketche of the Brillouin zone and the path of the integration in Eq. (15). The path $L$ connect high-symmetry points $\Gamma$, and $M$ ($M^{\prime}$). (f) Color plot of the $\mathbb{Z}_{4}$ Berry phase. (g) Schematic figures for two limits for $(\delta t,\gamma)=(1,1/2)$, $(\delta t,\gamma)\neq(\pm 1,\pm 1/2)$ and $(\delta t,\gamma)=(-1,-1/2)$.
@@
Two-Dimensional Su–Schrieffer–Heeger Model with the Nonreciprocal Hoppings: We consider the two-dimensional SSH model with the nonreciprocal hoppings [see Fig. 1 (a)]. The red and blue lines denote the nonreciprocal hoppings. Specifically, the hoppings along the red and blue arrows are $t_{1}+\gamma_{1}$ and $t_{2}+\gamma_{2}$, respectively. The hoppings opposite to the red and blue arrows are $t_{1}-\gamma_{1}$ and $t_{2}-\gamma_{2}$, respectively.
@@
${\footnotesize q_{2\times 2}(k_{1},k_{2})=\begin{pmatrix}0&t_{1}-\gamma_{1}+(t_{2}-\gamma_{2})e^{ik_{1}}\\ t_{1}+\gamma_{1}+(t_{2}+\gamma_{2})e^{ik_{2}}&0\end{pmatrix}.}$ (3)
@@
$\mathbb{Z}_{4}$ Berry Phase Protected by Generalized Four-Fold Rotational Symmetry: We see the topological invariant to characterize the real line gap topology. Specifically, we introduce the $\mathbb{Z}_{4}$ Berry phase protected by generalized four-fold symmetry in a two-dimensional momentum space. The key idea originates from the quantized Berry phase protected by the crystalline symmetry [16, 18, 24, 112, 113, 114, 116, 81]. In addition, the notion of the complex Berry phase also plays an important role [131, 132, 45, 35, 81]. Furthermore, in the presence of the crystalline symmetry, the form of the topological invariant can be simplified to one determined from the data at high-symmetry points in the Brillouin zone. This reduction is important from the view of the computational costs.
@@
To evaluate Eq. (15), we consider the constraints on the eigenvectors and the Berry connection arising from the presence of generalized four-fold symmetry. In the presence of generalized four-fold symmetry, the constraints on the right and left eigenvectors are
@@
The detailed derivation of Eq. (21) is presented in Supplemental Materials. We note that even if there are the exceptional points on the path $L$ except the high-symmetry points $\Gamma$ and $M$ in Fig. 1 (e), we can avoid them by the continuous deformation of the integration path. This is the unique advantage of employing the $\mathbb{Z}_{4}$ Berry phase, since the topological invariant such that the Chern number can not be defined when there exists an exceptional point in the two-dimensional Brillouin zone.
@@
In Fig.1 (f), we show the color map of $\mathbb{Z}_{4}$ phase of the band $E_{1}(\vec{k})$ in Eq. (21). The $\mathbb{Z}_{4}$ Berry phase takes values of $\text{Re}\hskip 1.0pt\nu=1/2$ (yellow area) and $\text{Re}\hskip 1.0pt\nu=0$ (blue area). The white area shows the existence of the exceptional rings around the high-symmetry point $M$.
@@
Fig. 2: (Color online) (a) Schematic figure for the boundary shape of the two-dimensional SSH model with nonreciprocal hoppings when considering the cylinder geometry. We arrange to break the inversion symmetry. The dashed line shows the center of the generalized mirror plane. (b)-(d) The red (green) dots show the energy spectrum under the periodic (open) boundary condition along the $x$ direction. for several parameters sets of $(\delta t,\delta\gamma)$. (b), (c), and (d) show the data for $(\delta t,\delta\gamma)=(-0.7,0.1)$, $(0,0.1)$, and $(0.7,0.1)$, respectively. (e) and (f) show the distribution of the right eigenvector in the real space of corner states for $(\delta t,\delta\gamma,E)=(-0.7,0.1,-1.51302+0.480218i)$ and $(0.7,0.1,-1.54824+0.260155i)$, respectively.
@@
Edge Spectrum on the Cylinder Geometry: Before going to the numerical result of the edge spectra, to get the insight of the relation between the $\mathbb{Z}_{4}$ Berry phase and edge states, we consider the two limits, i.e., $(\delta t,\delta\gamma)=(1,1/2)$ and $(\delta t,\delta\gamma)=(-1,-1/2)$. For $(\delta t,\delta\gamma)=(1,1/2)$, there exist dimers on the right and bottom sides, while there exist dimers on the left and top sides for $(\delta t,\delta\gamma)=(-1,-1/2)$. By combining with the discussion of the adiabatic connection of the $\mathbb{Z}_{4}$ Berry phase, $\text{Re}\hskip 1.0pt\nu=1/2$ (00) predicts the emergence of the edge states at left and top (right and bottom) sides. Thus, the $\mathbb{Z}_{4}$ Berry phase indicates the position of edge states.
@@
Once the profile of edge states are understood by the $\mathbb{Z}_{4}$ Berry phase, we can then argue the characterization of the skin effect of the edge spectra. To be more concrete, the non-zero winding number of the edge spectrum provides the emergence of the hybrid higher-order topological skin modes. The explicit form of the winding number of the edge spectra is
$w=\int^{\pi}_{-\pi}\frac{dk}{2\pi i}\hskip 1.0pt\frac{\partial}{\partial k}\log\det(h_{1\text{D}}(k)-E_{\text{ref}}),$ (26)
@@
where $h_{1\text{D}}(k)$ denotes the Hamiltonian on the cylinder geometry. However, in the presence of the conventional inversion symmetry [see Eq. (4)], the total of the winding number vanishes [75]. To avoid this problem, we consider the system in the absence of inversion symmetry by setting the appropriate edges [see Fig. 2 (a)]. The trivial case of edge spectra is discussed in Supplemental materials.
@@
We note that, in the present model, we can also define $\mathbb{Z}_{2}$ indicator with generalized mirror symmetry that that predicts the emergence of the skin effect and hybrid higher-order topological skin modes [75] [see Fig. 2 (a)]. The detail is discussed in Supplemental materials.
@@
Figures 2 (b), (c), and (d) display energy spectra under the periodic (open) boundary condition along $x$ ($y$) direction for $(\delta t,\delta\gamma)=(-0.7,0.1)$, $(0,0.1)$, and $(0.7,0.1)$, respectively. Here, we set the number of the unit cells as $20$. Thus, the total number of the lattice sites are $82$. There exists the energy spectrum of edge states where eigenvectors are localized at top and left [bottom and right] side for $(\delta t,\delta\gamma)=(-0.7,0.1)$ $[(0.7,0.1)]$. Here, we note that there may exist the energy spectrum of edge states even in the presence of gapless states. Thus, the emergence of all edge states are not predicted by the $\mathbb{Z}_{4}$ Berry phase.

### v2

# Hybrid Higher-Order Skin Topological Modes in the Two-Dimensional Su-Schrieffer-Heeger Model with Nonreciprocal Hoppings

@@ Abstract
The coexistence of edge states and skin effects provides the topologically protected localized states at the corners of two-dimensional systems. In this paper, we realize such corner states in the two-dimensional Su-Schrieffer-Heeger model with the nonreciprocal hoppings. For the characterization of the real line gap topology, we introduce the $\mathbb{Z}_{4}$ Berry phase protected by generalized four-fold rotational symmetry. From the physical picture of the adiabatic connection, we find that the value of the $\mathbb{Z}_{4}$ Berry phase predicts the position of edge states. Additionally, by using the winding number, we characterize the point gap topology of the edge spectra. From the results of these characterizations by the first-order topological invariants, we find that the pair of values of the $\mathbb{Z}_{4}$ Berry phase and the winding number yields the position of the topologically protected localized states.
1 Introduction
@@
Topological materials which are characterized by the topology in the bulk have been actively pursued [1, 2, 3, 4, 5, 6, 7, 8]. Additionally, the description of these materials are enriched by symmetry [9, 10, 11, 12, 13, 14, 15]. For instance, the presence of time-reversal symmetry provides $\mathbb{Z}_{2}$ topological properties [9]. Moreover, it turns out that crystalline symmetry plays an key role for the determination of the band topology [16, 17, 18, 19, 20, 21, 22, 23, 25, 24, 26]. For example, the coexistence of the time-reversal symmetry and inversion symmetry enables us to introduce the form of the $\mathbb{Z}_{2}$ indicator determined from the parity at the high-symmetry points in the two-dimensional Brillouin zone [17].
@@
In the past decades, the notion of the topological protection has been extended to a new area: the non-Hermitian systems [27, 28, 29, 30, 31, 32]. Since the eigenvalues may become complex, many exotic topological phenomena have been provided [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], such as the skin effects [48, 49, 50, 51, 52, 53, 54, 55, 56, 57] and exceptional points (rings) [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. Particularly, a distinct difference from the case in Hermitian systems is the emergence of two types of the gap, i.e. a line gap and a point gap [45]. While the line gap topology can be understood by mapping onto the Hermitian cases, the nontrivial point gap topology provides novel phenomena without analogies in the Hermitian case, such as the anomalous sensitivity of the energy spectrum to the presence or absence of boundaries, i.e., the skin effect. In addition, as is the case for the Hermitian systems, these unique topological phenomena are enriched by symmetry [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. These phenomena have been reported in a wide range of systems such as open quantum systems [59, 86, 87], electrical circuits [88, 89, 90, 74, 91, 92], phononic systems [93, 94, 95], and photonic crystals [96, 97, 98, 99], and so on.
@@
Along with the above developments, a novel class of topological insulators has been extensively investigated, which is called higher-order topological insulators. In $d$-dimensional systems, while the conventional topological insulators host $(d-1)$-dimensional boundary states, higher-order topological insulators host $(d-2)$- or fewer dimensional boundary states [100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 114, 113, 115, 116, 117, 118]. For the characterization of the topological phase and the prediction of the emergence of these unique boundary states, many efforts have been devoted. Remarkably, the presence of crystalline symmetry plays an important role for the description of the higher-order topology [100, 101, 102, 103, 104, 108, 109, 110, 111, 112, 114, 113, 115, 116, 117, 118].
@@
Recently, the notion of the higher-order topology has also been extended to non-Hermitian systems [119, 120, 121, 122]. In particular, the coexistence of edge states and the skin effect in two-dimensional systems provides the topologically protected corner states, which are called the hybrid higher-order skin topological modes [123, 124, 125, 126, 127, 128, 129]. These novel states in non-Hermitian systems represent a kind of the higher-order skin effect [130, 131, 132]. Furthermore, these unique localized states are characterized by the first- and second-order topological invariants such as the Chern number [123, 126], the Wess-Zumino term [130], and the symmetry indicator [131, 77].
@@
In this paper, we demonstrate the realization of the hybrid higher-order skin topological modes in the two-dimensional Su-Schrieffer-Heeger (SSH) model with the nonreciprocal hoppings. Specifically, we characterize the first-order topology by using the $\mathbb{Z}_{4}$ Berry phase protected by generalized four-fold rotational symmetry and the winding number on the cylinder geometry. From the nontrivial value of the $\mathbb{Z}_{4}$ Berry phase, we show that there exists the bulk-edge correspondence between this topological invariant and edge states. Additionally, by using the winding number, we realize the skin effect arising from the energy spectra of edge states. Combining the $\mathbb{Z}_{4}$ Berry phase and the nontrivial point gap topology of these edge spectra, we demonstrate that the pair of these topological invariants indicates the position of the localized corner in two dimension.
@@
Fig. 1: (Color online) (a) Schematic figure of the two-dimensional SSH model with the nonreciprocal hoppings. The unit cell is enclosed by the black dashed box. The cross denotes the center of the generalized four-fold rotation. (b), (c), and (d) show the energy spectrum for $(\delta t,\delta\gamma)=(-0.7,0.1)$, $(0,0.1)$ and $(0.7,0.1)$. Red lines denote the real line gaps where we focus. (e) Sketch of the Brillouin zone and the path of the integration in Eq. (15). The path $L$ connects high-symmetry points $\Gamma$ and $M$ ($M^{\prime}$). (f) Color plot of the $\mathbb{Z}_{4}$ Berry phase. (g) Schematic figures for $(\delta t,\gamma)=(1,1/2)$, $(\delta t,\gamma)\neq(\pm 1,\pm 1/2)$ and $(\delta t,\gamma)=(-1,-1/2)$.
The rest of this paper is organized as follows. In Sec. 2, we introduce the two-dimensional SSH model with nonreciprocal hoppings. In addition, we show the energy spectra under the periodic boundary condition. In Sec. 3, for the characterization of the real line gap topology, we first introduce the $\mathbb{Z}_{4}$ Berry phase protected by generalized four-fold rotational symmetry. We then show the explicit computation of the $\mathbb{Z}_{4}$ Berry phase in the present model. In Sec. 4, from the physical picture of the adiabatic connection, we discuss the emergence of the edge states on the cylinder geometry. Additionally, we characterize the point gap topology of edge spectra by the winding number. In Sec. 5, from the nontrivial point gap topology of the edge spectra, we show the emergence of the corner states characterized by the pair of the $\mathbb{Z}_{4}$ Berry phase and the winding number. In Sec. 6, we present a summery of this paper.
2 Two-dimensional Su-Schrieffer-Heeger model with the nonreciprocal hoppings
@@
We consider the two-dimensional SSH model with the nonreciprocal hoppings [see Fig. 1 (a)]. The red and blue lines denote the nonreciprocal hoppings. Specifically, the hoppings along the red and blue arrows are $t_{1}+\gamma_{1}$ and $t_{2}+\gamma_{2}$, respectively. The hoppings opposite to the red and blue arrows are $t_{1}-\gamma_{1}$ and $t_{2}-\gamma_{2}$, respectively.
@@
$q_{2\times 2}(k_{1},k_{2})=\begin{pmatrix}0&t_{1}-\gamma_{1}+(t_{2}-\gamma_{2})e^{ik_{1}}\\ t_{1}+\gamma_{1}+(t_{2}+\gamma_{2})e^{ik_{2}}&0\end{pmatrix}.$ (3)
@@
3 $\mathbb{Z}_{4}$ Berry phase protected by generalized four-fold rotational symmetry
We see the topological invariant to characterize the real line gap topology. Specifically, we introduce the $\mathbb{Z}_{4}$ Berry phase protected by generalized four-fold symmetry in a two-dimensional momentum space. The key idea originates from the quantized Berry phase protected by the crystalline symmetry [16, 18, 24, 113, 114, 115, 117, 118, 81]. Indeed, even in the discussion of the topology in the two-dimensional systems, the presence of rotational symmetry enables us to introduce the quantized Berry phase [18, 24, 114, 113, 115, 117, 118]. In addition, the notion of the complex Berry phase also plays an important role [133, 134, 46, 35, 81]. Furthermore, in the presence of the crystalline symmetry, the form of the topological invariant can be simplified to one determined from the data at high-symmetry points in the Brillouin zone. This reduction is important from the view of the computational costs.
@@ 3 $\mathbb{Z}_{4}$ Berry phase protected by generalized four-fold rotational symmetry
To evaluate Eq. (15), we consider the constraints on the eigenvectors and the Berry connection arising from the presence of generalized four-fold symmetry. In the presence of generalized four-fold symmetry, the constraints on right and left eigenvectors are
@@
The technical details regarding Eqs. (20) and (21) are shown in Appendixes A, B, and C. Even in the presence of gapless points on the integration path $L$ except the high-symmetry points $\Gamma$ and $M$ in Fig. 1 (e), by the continuous deformation of the integration path, we can employ the $\mathbb{Z}_{4}$ Berry phase for the characterization of the real line gap. We show the detail and the example in Appendix D. This is the unique advantage of employing the $\mathbb{Z}_{4}$ Berry phase, since the topological invariant such as the Chern number can not be defined when there exists an exceptional point in the two-dimensional Brillouin zone.
@@
In Fig.1 (f), we show the color map of the $\mathbb{Z}_{4}$ phase of the band $E_{1}(\vec{k})$ in Eq. (21). The $\mathbb{Z}_{4}$ Berry phase takes values of $\text{Re}\hskip 1.0pt\nu=1/2$ (yellow area) and $\text{Re}\hskip 1.0pt\nu=0$ (blue area). The white area shows the existence of the exceptional rings around the high-symmetry point $M$. We note that, in the present model, there are only two topological phases ($\text{Re}\hskip 1.0pt\nu=0$, $1/2$) whereas the Berry phase is quantized in $\mathbb{Z}_{4}$. In Appendix E, we show an example of the emergence of four topological phases in the different toy model.
@@
Fig. 2: (Color online) (a) Schematic figure for the boundary shape of the two-dimensional SSH model with nonreciprocal hoppings when considering the cylinder geometry. By breaking inversion symmetry, we avoid vanishing of the winding number. The dashed line shows the center of the generalized mirror plane. (b)-(d) The red (green) dots show the energy spectrum with the geometry in (a) under the periodic (open) boundary condition along the $x$ direction for several parameters sets of $(\delta t,\delta\gamma)$. (b), (c), and (d) show the data for $(\delta t,\delta\gamma)=(-0.7,0.1)$, $(0,0.1)$, and $(0.7,0.1)$, respectively. (e) Schematic illustration of the non-Hermitian two-dimensional SSH model with even number of sites along both $x$ and $y$ directions. (f)-(g) The red (green) dots show energy spectra under the periodic (open) boundary condition along the $x$ direction with the geometry in (e). (f) and (g) show the energy spectra on the complex plane for $(\delta t,\delta\gamma)=(-0.7,0.1)$ and $(0.7,0.1)$, respectively. (h) Schematic illustration of the non-Hermitian two-dimensional SSH model with even (odd) number of sites along the $x$ ($y$) direction. (i)-(j) The red (green) dots represent energy spectra under the periodic (open) boundary condition along the $x$ direction with the geometry in (h). (i) and (j) show the energy spectra on the complex plane for $(\delta t,\delta\gamma)=(-0.7,0.1)$, $(0.7,0.1)$, respectively.
4 Edge spectrum on the cylinder geometry
@@
Before going to the numerical result of the edge spectra, to get the insight of the relation between the $\mathbb{Z}_{4}$ Berry phase and edge states, we consider edge states originating from the dimers at the two limits, i.e., $(\delta t,\delta\gamma)=(1,1/2)$ and $(\delta t,\delta\gamma)=(-1,-1/2)$. Indeed, as shown in the previous work [113, 115], considering the adiabatic connection is a key to explain the correspondence between the Berry phase and the boundary modes, since it gives the simple real-space picture of the state. By the analogous approach, we consider the two certain limits in the present model [see Fig. 1 (g)]. Specifically, for $(\delta t,\delta\gamma)=(1,1/2)$, there exist dimers originating from the cut of the unit cell on the right and bottom sides, while there exist dimers on the left and top sides for $(\delta t,\delta\gamma)=(-1,-1/2)$. These dimers result in the emergence of edge states. By combining with the discussion of the adiabatic connection of the $\mathbb{Z}_{4}$ Berry phase, $\text{Re}\hskip 1.0pt\nu=1/2$ (00) predicts the emergence of the edge states at left and top (right and bottom) sides. Thus, the $\mathbb{Z}_{4}$ Berry phase indicates the position of edge states [135].
@@
Once the profile of edge states are understood by the $\mathbb{Z}_{4}$ Berry phase, we can then argue the characterization of the skin effect of the edge spectra. To be more concrete, the non-zero winding number of the edge spectrum provides the emergence of the hybrid higher-order skin topological modes. The explicit form of the winding number of the edge spectra is
$w=\int^{\pi}_{-\pi}\frac{dk}{2\pi i}\hskip 1.0pt\frac{\partial}{\partial k}\ln\det(h_{1\text{D}}(k)-E_{\text{ref}}1_{N\times N}),$ (26)
where $h_{1\text{D}}(k)$ and $E_{\text{ref}}$ denote the Hamiltonian on the cylinder geometry and the reference energy, respectively. However, in the presence of the conventional inversion symmetry [see Eq. (4)], the total of the winding number vanishes [75]. Specifically, the one-dimensional Hamiltonian preserves conventional inversion symmetry [75],
$U_{I}h_{1\text{D}}(k)U_{I}^{-1}=h_{1\text{D}}(-k),$ (27)
with $U_{I}$ being the unitary matrix ($U_{I}^{2}=1_{N\times N}$). The presence of this symmetry enables us to evaluate the winding number,
$\begin{split}w&=\int^{\pi}_{-\pi}\frac{dk}{2\pi i}\frac{\partial}{\partial k}\ln\det(h_{1\text{D}}(k)-E_{\text{ref}}1_{N\times N})\\ &=\int^{\pi}_{-\pi}\frac{dk}{2\pi i}\frac{\partial}{\partial k}\ln\det(U_{I}(h_{1\text{D}}(-k)-E_{\text{ref}}1_{N\times N})U_{I}^{-1})\\ &=-\int^{\pi}_{-\pi}\frac{dk}{2\pi i}\frac{\partial}{\partial k}\ln\det(h_{1\text{D}}(k)-E_{\text{ref}1_{N\times N}})\\ &=-w.\end{split}$ (28)
@@
Here, we have used Eq. (27) and the relation $\det(AB)=\det(BA)$. Thus, in the presence of the inversion symmetry, the total of the winding number vanishes. To avoid this problem, we consider the system in the absence of inversion symmetry by setting the appropriate edges [see Fig. 2 (a)]. The trivial case of edge spectra is discussed in Sec. 5.
Fig. 3: (Color online) (a) and (b) show the distribution of the right eigenvector in the real space of corner states with the geometry in Fig. 2 (a) for $(\delta t,\delta\gamma,E)=(-0.7,0.1,-1.51302+0.480218i)$ and $(0.7,0.1,-1.54824+0.260155i)$, respectively. (c) shows the distribution of the eigenstate with the geometry in Fig. 2 (e) for $(\delta t,\delta\gamma,E)=(-0.7,0.1,-1.50723526-0.47764771i)$. (d) and (e) show the distributions of the eigenstates with the geometry in Fig. 2 (h) for $(\delta t,\delta\gamma,E)=(-0.7,0.1,-1.52665387-0.52222989i)$ and $(0.7,0.1,-1.5395-0.193811i)$, respectively.
@@
We note that, in the present model, we can also define the $\mathbb{Z}_{2}$ indicator with generalized mirror symmetry that predicts the emergence of the skin effect [75] [see Fig. 2 (a)]. In the presence model, the Hamiltonian preserves the generalized mirror symmetry, namely, it satisfies
$U_{M}h_{1\text{D}}(k)U_{M}^{-1}=h^{\dagger}_{1\text{D}}(-k),$ (29)
where $U_{M}$ denotes the unitary operator. The presence of this symmetry enables us to define the $\mathbb{Z}_{2}$ indicator [75],
$\mu:=(-1)^{w}=\prod_{k=0,\pi}\text{sgn}(\det(h_{1\text{D}}(k)-E_{\text{ref}}1_{N\times N})).$ (30)
Here, we have assumed that the reference energy $E_{\text{ref}}$ is real. The nontrivial value of the indicator (i.e., $\mu=-1$) predicts the emergence of the skin effect, since it guarantees that $w$ is nonzeo. In the present system, the pair of the $\mathbb{Z}_{4}$ Berry phase and this $\mathbb{Z}_{2}$ indicator predicts the emergence of the hybrid higher-order skin topological modes. However, $\mu$ cannot predict the position of the hybrid higher-order skin topological modes since we cannot see the difference between $w=1$ and $w=-1$.
@@
Figures 2 (b), (c), and (d) display energy spectra under the periodic (open) boundary condition along $x$ ($y$) direction for $(\delta t,\delta\gamma)=(-0.7,0.1)$, $(0,0.1)$, and $(0.7,0.1)$, respectively. Here, we set the number of the unit cells as $20$. Thus, the total number of the lattice sites are $82$. There exists the energy spectrum of edge states where eigenvectors are localized at top and left [bottom and right] side for $(\delta t,\delta\gamma)=(-0.7,0.1)$ $[(0.7,0.1)]$. Here, we note that there may exist the energy spectrum of edge states even in the presence of gapless states. Thus, the emergence of all edge states is not predicted by the $\mathbb{Z}_{4}$ Berry phase.
Before closing this section, we see energy spectra in the presence of conventional inversion symmetry. Specifically, we consider the system with conventional inversion symmetry by setting appropriate boundaries [see Fig. 2 (e)]. Figure 2 (f) displays energy spectra under the periodic (open) boundary condition along the $x$ ($y$) direction for $(\delta t,\delta\gamma)=(-0.7,0.1)$. At first sight, this numerical result of edge spectra looks similar to the one in Fig. 2 (b). However, while there are two edge spectra in the absence of conventional inversion symmetry [see Fig. 2 (b)], there are four edge spectra in the presence of conventional inversion symmetry [see Fig. 2 (f)], each of which is doubly degenerated. In the presence of the inversion symmetry, one of edge spectra at the left side in Fig. 2 (f) is with $w=1$. Another one at the left side in Fig. 2 (f) is with $w=-1$. Then, the total of the winding number with conventional inversion symmetry vanishes. Therefore, in the conventional inversion symmetry, we cannot distinguish this case from the trivial loop of edge spectra. However, the corner states emerge even in the case for $w=0$. The detail is discussed in Sec. 5.
5 Localized states at the corner